# What should a neuron aim for? Designing local objective functions based on information theory

### v4

@@ 1 Introduction
Neuroscience models have indeed shown that local learning rules (Hebb, 1949, Dan and Poo, 2004, Song et al., 2000) can solve a variety of tasks, such as sequence learning, unsupervised clustering, and classification tasks (Hebb, 1949, Földiak, 1990, Cramer et al., 2020).
Often formulated on the basis of temporal coincidence (spike-timing), these very mechanistic rules do not offer direct insights into the actual information processing at each neuron. To facilitate such functional insight, more recently information theory has been used to formulate more abstract, but interpretable frameworks for local learning rules (e.g., Kay 2000, Wibral et al. 2017).
Information theory (Shannon, 1948) allows one to abstractly prescribe how much of the information from different inputs should be conveyed to the output of a neuron (Kay, 2000).
Moreover, it provides a general and flexible mathematical framework, which captures the information processing at an abstract and interpretable level free of details of the exact implementation. As such, information theory has been employed to define general optimization goals, including global objectives like cross-entropy loss (Goodfellow et al., 2016) and more localized ones such as local greedy infomax (Löwe et al., 2019), and early attempts at passing only information that is coherent between multiple inputs (Kay and Phillips, 2011).
@@
Next, we give a succinct account of infomorphic neurons with two input classes following Makkeh et al. (2025) before generalizing the framework to the more capable version with three input classes in Section 3.
@@ 2 Bivariate Infomorphic Neurons
Mutual information can be used to quantify the amount of information that is carried by different input classes about the neuron’s output by considering the aggregated feedforward ($F$), contextual ($C$) and lateral inputs ($L$) as sources $X$ individually. Nevertheless, classical information theory cannot quantify how much of the information is provided redundantly, uniquely or synergistically by different inputs, making it impossible to describe how exactly the information from two source variables contribute to creating the output. Using PID, the total mutual information $I(Y:F,C)$ between the output of the neuron $Y$ and two aggregated inputs, e.g., $F$ and $C$, can be dissected into four PID atoms (see Fig. 1.C): The unique information of the feedforward connection about the output $Y$, $\Pi_{\mathrm{unq,F}}$, is the information which can only be obtained from the feedforward and not the context input, with the unique information $\Pi_{\mathrm{unq,C}}$ of the context being defined analogously. The redundant information $\Pi_{\mathrm{red}}$ reflects the information which can equivalently obtained from either feedforward or contextual inputs about $Y$, while finally the synergistic information $\Pi_{\mathrm{syn}}$ can only be obtained from both inputs considered jointly. All classical mutual information terms between the target and subsets of source variables can be constructed from these PID atoms through the consistency equations (Williams and Beer, 2010)
@@
However, these atoms are underdetermined as there are four unknown atoms with only three consistency equations providing constraints. For this reason, an additional quantity needs to be defined, which is usually a measure for redundancy (Williams and Beer, 2010). Such a definition for redundancy, based on the concept of shared exclusions in probability space, is given by the shared-exclusion PID, $I^{\mathrm{sx}}_{\cap}$ (Makkeh et al., 2021) (see Section A.2). Importantly, the resulting PID is differentiable with respect to the underlying probability mass function, which allows for a specification and subsequent optimization of learning goals based on PID.
@@ 3 Trivariate Infomorphic Neurons
Figure 2: Adding lateral connections as a third input class enables neurons to self-organize to encode relevant and unique information. A,B. Infomorphic neuron with three inputs - namely feed-forward ($F$), contextual ($C$) and lateral ($L$) inputs. C. With three input classes, the number of PID-atoms increases to 18, represented by different colors, plus the residual entropy $H_{\mathrm{res}}$ in the outer circle. Classical information-theoretic quantities such as the entropy $H(Y)$ and mutual information $I$ with individual sources are depicted by ovals, indicating how they can be built from PID atoms. D. Three input classes allow for the optimization of more complex goals based on 19 different terms (compared to only five in Fig. 1.C). This trivariate PID allows to combine two bivariate objective functions: In a supervised learning task, the intuitive goal is to maximize information in the neuron’s output that is redundant between the feedforward input $F$ and label $C$, while simultaneously ensuring the neuron’s output stays unique with respect to lateral neurons $L$. While bivariate goal functions only allow for optimizing one of these objectives at a time, both objectives can be combined to the goal of maximizing the single atom $\Pi_{\{F\}\{C\}}$ in the trivariate case.
Including a third source variable results in a so-called trivariate PID which is more complex than the bivariate case, with atoms in general now describing redundancies between synergies. For example, the atom $\Pi_{\{F\}\{C\}}$ denotes information that can be obtained redundantly from the feedforward and context signal but is unavailable from the lateral inputs, making it ‘uniquely redundant’ to $F$ and $C$. Likewise, the atom $\Pi_{\{F\}\{C,L\}}$ denotes information that can be obtained from $F$ alone but also synergistically from the pair $C,L$ (but not from $C$ or $L$ alone). Overall, upon the introduction of a third source, the four PID atoms of the bivariate setup turn into a total of 18 trivariate PID atoms (Gutknecht et al., 2021) as shown in Fig. 2.C. These PID atoms $\Pi_{\alpha}$ can in general be addressed by their antichains $\alpha$, i.e., sets of sets of input variables, with the inner sets describing synergies and the outer set redundancies between them (Gutknecht et al., 2021).
@@
$G$ $=\gamma_{\{F\}\{C\}\{L\}}\Pi_{\{F\}\{C\}\{L\}}+...+\gamma_{\{FCL\}}\Pi_{\{F,C,L\}}+\gamma_{\mathrm{res}}H_{\mathrm{res}}=\bm{\gamma}^{T}\mathbf{\Pi}.$ (3)
Utilizing the expressiveness and interpretability of trivariate PID, heuristic goal functions can be derived for different classes of tasks from intuitive reasoning. For the supervised learning task to be solved here, the introduction of the lateral connections as a third input allows to synthesize an intuitive goal function from the two bivariate goals introduced before (see Fig. 2.D). As outlined in the bivariate case above, canonical goals might either be to foster redundancy between the feedforward and context signals with no regards for what lateral neurons do, or, alternatively, to optimize for uniqueness of the output with respect to the lateral inputs, which does not differentiate between task relevance of information. With trivariate infomorphic neurons, however, these goals can simultaneously be achieved by optimizing for the atom $\Pi_{\{F\}\{C\}}$, which represents the redundancy of the feedforward and context signal which is at the same time unique with respect to the lateral inputs. In simple terms, this enables the neuron to find coherent information between label and input data, that is not yet encoded by other neurons.
@@ 4 Experiments
Neuron structure and activation function   To construct infomorphic networks, a number of practical decisions need to be made for how the output signal is constructed from the inputs in the forward pass. As a first step, the higher-dimensional signals from the different input classes are aggregated to the single numbers $F$, $C$ and $L$ by a weighted sum. Subsequently, these aggregated inputs are passed into an activation function, which can be chosen arbitrarily, but needs to fit the requirements of the specific application. For the supervised classification task at hand, the function $A(F,C,L)=F[(1-\alpha_{1}-\alpha_{2})+\alpha_{1}\,\sigma(\beta_{1}FC)+\alpha_{2}\,\sigma(\beta_{2}FL)]$ has been chosen, where $\sigma$ refers to the sigmoid function and $\alpha_{1}=\alpha_{2}=0.1$ and $\beta_{1}=\beta_{2}=2$ are parameters that shape the influence of the input compartments on the activation function. This activation function builds on concepts of Kay (1994) and makes a clear distinction between the driving feedforward input and the modulatory context and lateral inputs, ensuring that the network performs similarly during training, when the context inputs is provided, and for evaluation, where it is withheld (i.e., by setting $C=0$). To show the independence of the infomorphic approach from a particular activation function, the results are additionally compared to a simple linear activation function $A(F,C,L)=F+0.1L+0.1C$. The output of the activation function is finally mapped to the unit interval by a sigmoid function, whose output is used as a probability to stochastically output “1” (firing) or “-1” (non-firing). For the bivariate output layer, the sigmoid is not sampled from but just interpreted as a firing probability, with the neuron with the highest value being used as the network prediction. The lateral connections between neurons add recurrence to the network, requiring for the same input to be shown multiple times to ensure convergence of the activations. In practice, however, we have seen that presenting the same input twice is sufficient to achieve proper network function (see Section A.9). The inference procedure for infomorphic networks is summarized in pseudocode in Section A.1.
@@
Discretization and local gradients   To optimize a PID-based goal function during training, the first step is to estimate the PID atoms. Since the original $I^{\mathrm{sx}}_{\cap}$ measure only works with discrete variables, the aggregated inputs from a batch of samples are first discretized to $20$ levels each (see Section A.11). Subsequently, the empirical probability mass function is constructed from the batch, from which the PID atoms are computed by means of a plug-in estimation. Finally, the current value of the goal function can determined as a linear combination of PID atoms (Algorithm 2 and Algorithm 3 in Section A.1). For the local gradient computation, the gradient of the individual neuron’s objective function is backpropagated via the output to produce the neuron’s weight updates.
@@
Network Setup   To solve the MNIST task (LeCun et al., 1998), we devised a network architecture consisting of an input layer, a hidden layer made up from $N_{\mathrm{hid}}$ trivariate infomorphic neurons and an output layer consisting of $10$ bivariate infomorphic neurons. The trivariate neurons in the hidden layer each receive the whole image as their feedforward signal, a context signal that guides their learning, and the output of other neurons of the same layer through lateral connections. The bivariate neurons in the output layer receive all outputs of the hidden layer and additionally exactly one element of the one-hot label as context, which is only provided during training, and are trained to maximize redundancy between their input and the information about the one label they observe. We explored three slightly different approaches to how the hidden layer can be set up using trivariate infomorphic neurons, illustrated in Fig. 3.A. In setup 1, we use the fully connected one-hot label as $C$ and the output of all neurons of the same layer as all-to-all connected lateral input $L$. In setup 2, a sparse connectivity of a maximum of 100 connections is used instead of all-to-all for $L$. In setup 3, additionally to sparse $L$, the label is replaced by a fully connected feedback from the output layer as context $C$, indicating a path towards how multiple hidden layers may be stacked in the future. For comparison, we trained a benchmark network with the same connectivity as setup 1 using standard backpropagation and cross-entropy loss. Additionally, we trained the output layer of a fixed random hidden layer setup using step-function activation.
@@ 5 Results
Figure 3: Infomorphic networks with three input classes approach the performance of a similar network trained with backpropagation and outperform networks with two input classes on the MNIST handwritten digit classification task. A. In the hidden layer of the network, the infomorphic neurons receive feedforward and lateral connections as well as either the ground-truth label or feedback from the output layer, outlined in three setups. B. Networks with trivariate infomorphic neurons outperform models with bivariate neurons with both, heuristic and optimized goals for a hidden layer size of 100 neurons. Only the setup with a feedback signal instead of the label performs worse for this layer size but outperforms the bivariate models for larger hidden layers as shown in Fig. 8. C. Infomorphic networks achieve similar performance as the same network trained with backpropagation. For larger layers, using a sparse connectivity significantly improves performance (setup 2). The lines indicate mean values, with the intervals depicting the maximum and minimum of 10 runs. The goal function parameters have been optimized for networks with a hidden layers size of 100 neurons, indicated by the dashed line in C.
@@
Performance   The performance of all major setups for $N_{\mathrm{hid}}=100$ is shown in Fig. 3.B. The three trivariate infomorphic setups use the same set of optimized goal parameters (see below) for all hidden layer sizes and significantly outperform the random baseline as well as the networks with a bivariate setup for the hidden layer. As shown in Appendix Fig. 3.C, Setup 1 matches the performance of backpropagation for up to 100 neurons, and reaches its maximum test accuracy for 500 neurons before its performance starts decreasing with larger layer sizes. This decrease in accuracy can be attributed to a lack of convergence of the neurons (Appendix Fig. 5), likely arising through the interaction of too many neurons. To alleviate the convergence problem, we performed additional experiments (setup 2) with the number of lateral connections reduced to maximally 100, which leads to better convergence and contributes to the continuous increase of performance for larger hidden layers, reaching a median test accuracy of $97.5\%$ for 2000 neurons, slightly below the $98.0\%$ reached via backpropagation. Finally, experiments with setup 3 provide evidence that the direct label input to the hidden layer can be replaced with feedback from the next layer, while still enabling solid performance especially for large layers. Additionally, we tested the alternate linear activation function as outlined in 4. Here, experiments using the optimized goal function parameters resulted in accuracies between $94.9\,\%$ and $95.1\,\%$ for $100$ neurons, which is on par with the original activation function.
@@
Figure 4: Given a set of optimized goal parameters, an ablation study allows for the identification and interpretation of the most critical neural subgoals for a given task. A. The heuristically defined goal function (see Fig. 2.D) shows 92.8% test accuracy on MNIST for $N_{hid}=100$; optimizing the goal function using hyperparameter optimization increases the test accuracy to 94.9%. The optimized goal parameters include the intuitive $\gamma_{\{F\}\{C\}}$, but also additional PID atoms to be maximized or minimized at the same time. B. For identifying the most important goal parameters, we performed an ablation study (individually setting $\gamma$ parameters to 0) and measured the change in validation accuracy compared to a network trained with the full goal. C. A successive ablation of parameters in order from lowest to highest individual effect identifies four parameters as being crucial for network performance (see Table 3 for their definition). The lines indicate mean values, with the intervals depicting the maximum and minimum of 10 runs. D. To test whether more complex image classification tasks require different goal parameters, we perform a separate hyperparameter optimization for setup 1, 100 neurons in CIFAR-10 and reach a median test accuracy of $42.5\%$ (compared to $42.2\%$ using backprop and $41.1\%$ using the goal function optimized for MNIST).
@@
Goal parameters   The optimized goal function used for training the three main setups at all sizes were obtained by maximizing the validation set accuracy using setup 1 with $N_{\mathrm{hid}}=100$. We performed a total of six hyperparameter optimizations of 1000 trials, three using TPE and three using CMA-ES samplers, respectively (see Appendix Fig. 6). The best results were achieved using the parameters shown in Fig. 4. The optimized objective functions outperform the intuitive goal function by including additional PID atoms besides the intuitively derived $\Pi_{\{F\}\{C\}}$. To compare to a more complex image classification task, a single CMA-ES optimization run has been performed for the CIFAR-10 task (Krizhevsky, 2012), with the optimal parameters presented in Fig. 4.D.
@@
Parameter importance   An analysis of the optimized hyperparameter values (Fig. 4.B) shows that only relatively few parameters are of high importance for performance. By setting goal parameters to zero individually (Figure 4.B) we find that there are four goal parameters that are critically non-zero for high performance on MNIST. This finding is confirmed by cumulatively setting goal function parameters to zero in the order of the individual drop in performance, which is illustrated in Fig. 4.C. Interestingly, we find that CIFAR-10 two of the four critical goal parameters for MNIST seem especially important. As a complementary measure of parameter importance, the results of mean decrease in impurity (Breiman, 2001) can be found in Section A.6. The results of a more detailed analysis of parameter importance are presented in Appendix Fig. 9.
@@
Interpretation   For the MNIST task, neurons aim for (i) encoding information that is redundant between feedforward and context inputs and thus task relevant, but not already encoded in other neurons (strongly positive $\gamma_{\{F\}\{C\}}$), (ii) avoid encoding information that is not in the context, thus not task-relevant, and already encoded by other neurons (strongly negative $\gamma_{\{F\}\{L\}}$) and (iii) allow redundancies of task relevant information between neurons (slightly positive $\gamma_{\{F\}\{C\}\{L\}}$). Additionally, neurons avoid synergies that require $F$ and either $C$ or $L$ for their recovery ($\gamma_{\{FC\}\{FL\}}$) (see Appendix Table 3 for an overview of the parameter values and interpretations). For the CIFAR-10 task (see Fig. 4.D), the two most important parameters $\gamma_{\{F\}\{C\}\{L\}}$ and $\gamma_{\{F\}\{C\}}$ point to a prioritization of redundancy over uniqueness between neurons.

### v5

@@ 1 Introduction
Neuroscience models have indeed shown that local learning rules (Hebb, 1949, Dan and Poo, 2004, Song et al., 2000) can solve a variety of tasks, such as sequence learning, unsupervised clustering, and classification tasks (Hebb, 1949, Földiak, 1990, Cramer et al., 2020). Often formulated on the basis of temporal coincidence (spike-timing), these very mechanistic rules do not offer direct insights into the actual information processing at each neuron. To facilitate such functional insight, more recently information theory has been used to formulate more abstract, but interpretable frameworks for local learning rules (e.g., Kay 2000, Wibral et al. 2017).
Information theory (Shannon, 1948) allows one to abstractly prescribe how much of the information from different inputs should be conveyed to the output of a neuron (Kay, 2000). Moreover, it provides a general and flexible mathematical framework, which captures the information processing at an abstract and interpretable level free of details of the exact implementation. As such, information theory has been employed to define general optimization goals, including global objectives like cross-entropy loss (Goodfellow et al., 2016) and more localized ones such as local greedy infomax (Löwe et al., 2019), and early attempts at passing only information that is coherent between multiple inputs (Kay and Phillips, 2011).
@@
Next, we give a succinct account of infomorphic neurons with two input classes following Makkeh et al. (2025) before generalizing the framework to the more capable version with three input classes in Section˜3.
@@ 2 Bivariate Infomorphic Neurons
Mutual information can be used to quantify the amount of information that is carried by different input classes about the neuron’s output by considering the aggregated feedforward ($F$), contextual ($C$) and lateral inputs ($L$) as sources $X$ individually. Nevertheless, classical information theory cannot quantify how much of the information is provided redundantly, uniquely or synergistically by different inputs, making it impossible to describe how exactly the information from two source variables contribute to creating the output. Using PID, the total mutual information $I(Y:F,C)$ between the output of the neuron $Y$ and two aggregated inputs, e.g., $F$ and $C$, can be dissected into four PID atoms (see Fig.˜1.C): The unique information of the feedforward connection about the output $Y$, $\Pi_{\mathrm{unq,F}}$, is the information which can only be obtained from the feedforward and not the context input, with the unique information $\Pi_{\mathrm{unq,C}}$ of the context being defined analogously. The redundant information $\Pi_{\mathrm{red}}$ reflects the information which can equivalently obtained from either feedforward or contextual inputs about $Y$, while finally the synergistic information $\Pi_{\mathrm{syn}}$ can only be obtained from both inputs considered jointly. All classical mutual information terms between the target and subsets of source variables can be constructed from these PID atoms through the consistency equations (Williams and Beer, 2010)
@@
However, these atoms are underdetermined as there are four unknown atoms with only three consistency equations providing constraints. For this reason, an additional quantity needs to be defined, which is usually a measure for redundancy (Williams and Beer, 2010). Such a definition for redundancy, based on the concept of shared exclusions in probability space, is given by the shared-exclusion PID, $I^{\mathrm{sx}}_{\cap}$ (Makkeh et al., 2021) (see Section˜A.2). Importantly, the resulting PID is differentiable with respect to the underlying probability mass function, which allows for a specification and subsequent optimization of learning goals based on PID.
@@ 3 Trivariate Infomorphic Neurons
Figure 2: Adding lateral connections as a third input class enables neurons to self-organize to encode relevant and unique information. A,B. Infomorphic neuron with three inputs - namely feed-forward ($F$), contextual ($C$) and lateral ($L$) inputs. C. With three input classes, the number of PID-atoms increases to 18, represented by different colors, plus the residual entropy $H_{\mathrm{res}}$ in the outer circle. Classical information-theoretic quantities such as the entropy $H(Y)$ and mutual information II with individual sources are depicted by ovals, indicating how they can be built from PID atoms. D. Three input classes allow for the optimization of more complex goals based on 19 different terms (compared to only five in Fig.˜1.C). This trivariate PID allows to combine two bivariate objective functions: In a supervised learning task, the intuitive goal is to maximize information in the neuron’s output that is redundant between the feedforward input $F$ and label $C$, while simultaneously ensuring the neuron’s output stays unique with respect to lateral neurons $L$. While bivariate goal functions only allow for optimizing one of these objectives at a time, both objectives can be combined to the goal of maximizing the single atom $\Pi_{\{F\}\{C\}}$ in the trivariate case.
Including a third source variable results in a so-called trivariate PID which is more complex than the bivariate case, with atoms in general now describing redundancies between synergies. For example, the atom $\Pi_{\{F\}\{C\}}$ denotes information that can be obtained redundantly from the feedforward and context signal but is unavailable from the lateral inputs, making it ‘uniquely redundant’ to $F$ and $C$. Likewise, the atom $\Pi_{\{F\}\{C,L\}}$ denotes information that can be obtained from $F$ alone but also synergistically from the pair $C,L$ (but not from $C$ or $L$ alone). Overall, upon the introduction of a third source, the four PID atoms of the bivariate setup turn into a total of 18 trivariate PID atoms (Gutknecht et al., 2021) as shown in Fig.˜2.C. These PID atoms $\Pi_{\alpha}$ can in general be addressed by their antichains $\alpha$, i.e., sets of sets of input variables, with the inner sets describing synergies and the outer set redundancies between them (Gutknecht et al., 2021).
@@
$G$ $=\gamma_{\{F\}\{C\}\{L\}}\Pi_{\{F\}\{C\}\{L\}}+.+\gamma_{\{FCL\}}\Pi_{\{F,C,L\}}+\gamma_{\mathrm{res}}H_{\mathrm{res}}=\bm{\gamma}^{T}\mathbf{\Pi}.$ (3)
Utilizing the expressiveness and interpretability of trivariate PID, heuristic goal functions can be derived for different classes of tasks from intuitive reasoning. For the supervised learning task to be solved here, the introduction of the lateral connections as a third input allows to synthesize an intuitive goal function from the two bivariate goals introduced before (see Fig.˜2.D). As outlined in the bivariate case above, canonical goals might either be to foster redundancy between the feedforward and context signals with no regards for what lateral neurons do, or, alternatively, to optimize for uniqueness of the output with respect to the lateral inputs, which does not differentiate between task relevance of information. With trivariate infomorphic neurons, however, these goals can simultaneously be achieved by optimizing for the atom $\Pi_{\{F\}\{C\}}$, which represents the redundancy of the feedforward and context signal which is at the same time unique with respect to the lateral inputs. In simple terms, this enables the neuron to find coherent information between label and input data, that is not yet encoded by other neurons.
@@ 4 Experiments
Neuron structure and activation function   To construct infomorphic networks, a number of practical decisions need to be made for how the output signal is constructed from the inputs in the forward pass. As a first step, the higher-dimensional signals from the different input classes are aggregated to the single numbers $F$, $C$ and $L$ by a weighted sum. Subsequently, these aggregated inputs are passed into an activation function, which can be chosen arbitrarily, but needs to fit the requirements of the specific application. For the supervised classification task at hand, the function $A(F,C,L)=F[(1-\alpha_{1}-\alpha_{2})+\alpha_{1}\,\sigma(\beta_{1}FC)+\alpha_{2}\,\sigma(\beta_{2}FL)]$ has been chosen, where $\sigma$ refers to the sigmoid function and $\alpha_{1}=\alpha_{2}=0.1$ and $\beta_{1}=\beta_{2}=2$ are parameters that shape the influence of the input compartments on the activation function. This activation function builds on concepts of Kay (1994) and makes a clear distinction between the driving feedforward input and the modulatory context and lateral inputs, ensuring that the network performs similarly during training, when the context inputs is provided, and for evaluation, where it is withheld (i.e., by setting $C=0$). To show the independence of the infomorphic approach from a particular activation function, the results are additionally compared to a simple linear activation function $A(F,C,L)=F+0.1L+0.1C$. The output of the activation function is finally mapped to the unit interval by a sigmoid function, whose output is used as a probability to stochastically output “1” (firing) or “-1” (non-firing). For the bivariate output layer, the sigmoid is not sampled from but just interpreted as a firing probability, with the neuron with the highest value being used as the network prediction. The lateral connections between neurons add recurrence to the network, requiring for the same input to be shown multiple times to ensure convergence of the activations. In practice, however, we have seen that presenting the same input twice is sufficient to achieve proper network function (see Section˜A.9). The inference procedure for infomorphic networks is summarized in pseudocode in Section˜A.1.
@@
Discretization and local gradients   To optimize a PID-based goal function during training, the first step is to estimate the PID atoms. Since the original $I^{\mathrm{sx}}_{\cap}$ measure only works with discrete variables, the aggregated inputs from a batch of samples are first discretized to $20$ levels each (see Section˜A.11). Subsequently, the empirical probability mass function is constructed from the batch, from which the PID atoms are computed by means of a plug-in estimation. Finally, the current value of the goal function can determined as a linear combination of PID atoms (Algorithm˜2 and Algorithm˜3 in Section˜A.1). For the local gradient computation, the gradient of the individual neuron’s objective function is backpropagated via the output to produce the neuron’s weight updates.
@@
Network Setup   To solve the MNIST task (LeCun et al., 1998), we devised a network architecture consisting of an input layer, a hidden layer made up from $N_{\mathrm{hid}}$ trivariate infomorphic neurons and an output layer consisting of $10$ bivariate infomorphic neurons. The trivariate neurons in the hidden layer each receive the whole image as their feedforward signal, a context signal that guides their learning, and the output of other neurons of the same layer through lateral connections. The bivariate neurons in the output layer receive all outputs of the hidden layer and additionally exactly one element of the one-hot label as context, which is only provided during training, and are trained to maximize redundancy between their input and the information about the one label they observe. We explored three slightly different approaches to how the hidden layer can be set up using trivariate infomorphic neurons, illustrated in Fig.˜3.A. In setup 1, we use the fully connected one-hot label as $C$ and the output of all neurons of the same layer as all-to-all connected lateral input $L$. In setup 2, a sparse connectivity of a maximum of 100 connections is used instead of all-to-all for $L$. In setup 3, additionally to sparse $L$, the label is replaced by a fully connected feedback from the output layer as context $C$, indicating a path towards how multiple hidden layers may be stacked in the future. For comparison, we trained a benchmark network with the same connectivity as setup 1 using standard backpropagation and cross-entropy loss. Additionally, we trained the output layer of a fixed random hidden layer setup using step-function activation.
@@ 5 Results
Figure 3: Infomorphic networks with three input classes approach the performance of a similar network trained with backpropagation and outperform networks with two input classes on the MNIST handwritten digit classification task. A. In the hidden layer of the network, the infomorphic neurons receive feedforward and lateral connections as well as either the ground-truth label or feedback from the output layer, outlined in three setups. B. Networks with trivariate infomorphic neurons outperform models with bivariate neurons with both, heuristic and optimized goals for a hidden layer size of 100 neurons. Only the setup with a feedback signal instead of the label performs worse for this layer size but outperforms the bivariate models for larger hidden layers as shown in Fig.˜8. C. Infomorphic networks achieve similar performance as the same network trained with backpropagation. For larger layers, using a sparse connectivity significantly improves performance (setup 2). The lines indicate mean values, with the intervals depicting the maximum and minimum of 10 runs. The goal function parameters have been optimized for networks with a hidden layers size of 100 neurons, indicated by the dashed line in C.
@@
Performance   The performance of all major setups for $N_{\mathrm{hid}}=100$ is shown in Fig.˜3.B. The three trivariate infomorphic setups use the same set of optimized goal parameters (see below) for all hidden layer sizes and significantly outperform the random baseline as well as the networks with a bivariate setup for the hidden layer. As shown in Appendix Fig.˜3.C, Setup 1 matches the performance of backpropagation for up to 100 neurons, and reaches its maximum test accuracy for 500 neurons before its performance starts decreasing with larger layer sizes. This decrease in accuracy can be attributed to a lack of convergence of the neurons (Appendix Fig.˜5), likely arising through the interaction of too many neurons. To alleviate the convergence problem, we performed additional experiments (setup 2) with the number of lateral connections reduced to maximally 100, which leads to better convergence and contributes to the continuous increase of performance for larger hidden layers, reaching a median test accuracy of $97.5\%$ for 2000 neurons, slightly below the $98.0\%$ reached via backpropagation. Finally, experiments with setup 3 provide evidence that the direct label input to the hidden layer can be replaced with feedback from the next layer, while still enabling solid performance especially for large layers. Additionally, we tested the alternate linear activation function as outlined in 4. Here, experiments using the optimized goal function parameters resulted in accuracies between $94.9\,\%$ and $95.1\,\%$ for $100$ neurons, which is on par with the original activation function.
@@
Figure 4: Given a set of optimized goal parameters, an ablation study allows for the identification and interpretation of the most critical neural subgoals for a given task. A. The heuristically defined goal function (see Fig.˜2.D) shows 92.8% test accuracy on MNIST for $N_{hid}=100$; optimizing the goal function using hyperparameter optimization increases the test accuracy to 94.9%. The optimized goal parameters include the intuitive $\gamma_{\{F\}\{C\}}$, but also additional PID atoms to be maximized or minimized at the same time. B. For identifying the most important goal parameters, we performed an ablation study (individually setting $\gamma$ parameters to 0) and measured the change in validation accuracy compared to a network trained with the full goal. C. A successive ablation of parameters in order from lowest to highest individual effect identifies four parameters as being crucial for network performance (see Table˜3 for their definition). The lines indicate mean values, with the intervals depicting the maximum and minimum of 10 runs. D. To test whether more complex image classification tasks require different goal parameters, we perform a separate hyperparameter optimization for setup 1, 100 neurons in CIFAR-10 and reach a median test accuracy of $42.5\%$ (compared to $42.2\%$ using backprop and $41.1\%$ using the goal function optimized for MNIST).
@@
Goal parameters   The optimized goal function used for training the three main setups at all sizes were obtained by maximizing the validation set accuracy using setup 1 with $N_{\mathrm{hid}}=100$. We performed a total of six hyperparameter optimizations of 1000 trials, three using TPE and three using CMA-ES samplers, respectively (see Appendix Fig.˜6). The best results were achieved using the parameters shown in Fig.˜4. The optimized objective functions outperform the intuitive goal function by including additional PID atoms besides the intuitively derived $\Pi_{\{F\}\{C\}}$. To compare to a more complex image classification task, a single CMA-ES optimization run has been performed for the CIFAR-10 task (Krizhevsky, 2012), with the optimal parameters presented in Fig.˜4.D.
@@
Parameter importance   An analysis of the optimized hyperparameter values (Fig.˜4.B) shows that only relatively few parameters are of high importance for performance. By setting goal parameters to zero individually (Figure 4.B) we find that there are four goal parameters that are critically non-zero for high performance on MNIST. This finding is confirmed by cumulatively setting goal function parameters to zero in the order of the individual drop in performance, which is illustrated in Fig.˜4.C. Interestingly, we find that CIFAR-10 two of the four critical goal parameters for MNIST seem especially important. As a complementary measure of parameter importance, the results of mean decrease in impurity (Breiman, 2001) can be found in Section˜A.6. The results of a more detailed analysis of parameter importance are presented in Appendix Fig.˜9.
@@
Interpretation   For the MNIST task, neurons aim for (i) encoding information that is redundant between feedforward and context inputs and thus task relevant, but not already encoded in other neurons (strongly positive $\gamma_{\{F\}\{C\}}$), (ii) avoid encoding information that is not in the context, thus not task-relevant, and already encoded by other neurons (strongly negative $\gamma_{\{F\}\{L\}}$) and (iii) allow redundancies of task relevant information between neurons (slightly positive $\gamma_{\{F\}\{C\}\{L\}}$). Additionally, neurons avoid synergies that require $F$ and either $C$ or $L$ for their recovery ($\gamma_{\{FC\}\{FL\}}$) (see Appendix Table˜3 for an overview of the parameter values and interpretations). For the CIFAR-10 task (see Fig.˜4.D), the two most important parameters $\gamma_{\{F\}\{C\}\{L\}}$ and $\gamma_{\{F\}\{C\}}$ point to a prioritization of redundancy over uniqueness between neurons.